OpenStudy (anonymous):

What mass of ethylene glycol (C2H6O2, molecular weight 62.1 g) must be added to 20.00 L of water to produce a solution for use in a car's radiator that freezes at -15 degrees centigrade? (The density of water is 1 Kg L^-1. The molal freezing point depression constant for water, Kf, is 1.86 degrees centigrade Kg mol^-1)

OpenStudy (anonymous):

delta T = Kf m delta T = 15K Kf= 1.86 Let x be the mass of ethylene glycol, then number of moles of ehtylene glycol will be x/62.1 Molality m = x/(62.1*20) So, 15=1.86*x/(62.1*20) x=15*62.1*20/1.86 = 10016.129 gram =10 kg approx